OpenStudy (andijo76):

use the given data set to x=10, 8, 13,9,11,14,6,4,12,7,5 y=9.15,8.14,8.75,8.76,9.25,8.11,6.13,3.09,9.13,7.26,4.73

OpenStudy (anonymous):

draw a graph?

OpenStudy (andijo76):

find the linear correlation coefficient r then determine whether there is sufficient evidence to support the claim of a linear correlation between two variables

OpenStudy (andijo76):

x=10,18,13,9,11,14,6,4,12,7,5 y=9.15,8.14,8.76,9.25,8.11,6.13,7.26,4.73 x*y=91.5,65.12,113.75,78.84,101.75,113.54,36.78,12.36,109.56,50.82 x^2=100,64,169,81,121,196,36,16,144,49,25 y^2=83.72,66.72,76.56,76.74,85.56,65.77,37.58,9.55,83.36,52.71,22.37 n=11 x=99 y=82.5 x^2=1001 y^2=651.18 this is what i came up with so far

OpenStudy (amistre64):

using average x and average y for the "point" of a line and a slope defined by: \[m=\frac{n\sum xy-\sum x \sum y}{n\sum xx - \sum x \sum x}\] there is some way thats a bit foggy at the moment to determine r

OpenStudy (amistre64):

there is some covariance formula that starts out like this: sum (xi-x')(yi-y') = sum (xiyi -xiy' -yix' +x'y')

OpenStudy (amistre64):

ti83 is soo much simpler .... or an excel run